OpenStudy (anonymous):

Let X =Y = { x in R: 0 < x<1 } Define h: XxY -> R by h(x,y) = 2x + y a. for each x in X find f(x) = sup{h(x,y): y in Y} then find inf {f(x): x in X} b. for each y find g(y) = inf {h(x,y): x in X} then find sup {g(y): y in Y}.

OpenStudy (jamesj):

Ok, so, is f(x) clear? Can you write it down?

OpenStudy (jamesj):

f(x) = sup{ 2x + y | y in (0,1) } hence f(x) = ....

OpenStudy (anonymous):

yes that is f(x) it is thwe result when you map x,y onto h(x)

OpenStudy (jamesj):

I want you to see there is way to write a formula for f(x) that doesn't involve sup or inf. Once you have formula, we can then deal with the second part of inf{f(x) : ....}

OpenStudy (jamesj):

If you're not seeing this, let me ask you this. What is sup{ y | y in (0,1) } ?

OpenStudy (anonymous):

yes that is f(x) it is thwe result when you map x,y onto h(x)

OpenStudy (anonymous):

f(x) = sup (2x +y) = 2x + 1 since y 0<y<1

OpenStudy (jamesj):

Right. Hence inf{ f(x) | x in (0,1) } is what?

OpenStudy (anonymous):

yes that is f(x) it is thwe result when you map x,y onto h(x)

OpenStudy (anonymous):

0

OpenStudy (jamesj):

Nope

OpenStudy (jamesj):

inf{ f(x) | x in (0,1) } = inf{ 2x + 1 | x in (0,1) } = ...

OpenStudy (anonymous):

yes that is f(x) it is thwe result when you map x,y onto h(x)

OpenStudy (anonymous):

oops of f(x) = inf (2x +1 ) = 1

OpenStudy (jamesj):

right. I'll you work through the second part.

OpenStudy (jamesj):

I'll let you work through the second part.

OpenStudy (anonymous):

yes that is f(x) it is thwe result when you map x,y onto h(x)

OpenStudy (anonymous):

g(y) = y

OpenStudy (anonymous):

yes that is f(x) it is thwe result when you map x,y onto h(x)

OpenStudy (anonymous):

sup(g(y)) = 1

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

yes that is f(x) it is thwe result when you map x,y onto h(x)

OpenStudy (anonymous):

coolio and beans. thanks. now I have to do it in another problem for h (x,y) = 0 if x << y or 1 if x[\le\] y. just vary the x for f(x) or y for g(y) and just follow through as above?

OpenStudy (jamesj):

Put this in a new post and make sure you have a well posed/defined question, as I'm not following exactly what you're saying.

OpenStudy (anonymous):

yes that is f(x) it is thwe result when you map x,y onto h(x)

OpenStudy (anonymous):

regardless of what function h(x,y) is, would the reasoning to find inf (f(x)) and sup g(y) the same?

OpenStudy (jamesj):

In general, yes.

OpenStudy (anonymous):

yes that is f(x) it is thwe result when you map x,y onto h(x)

OpenStudy (anonymous):

thanks again for the walk through. if you can throw any hints at the other posted question that would be cool.